A 30-year bond with par interest. The amount for accumulation of discount in the tenth coupon is 5.7963. Calculate the value 1000 and 6% annual coupons is purchased to yield 7% annual effective redemption value.
Answer #1

Solution :

F = 1000, r = 6% , i = 7% , v = 1 /( 1 +yield) = 1 / 1.07 , accumulation of discount = 5.7963 in 10th period ,

n =30 , t =10 , C =?

In order to find redemption value, we can use the bond amortization formula

(Accumulation of discount)t = (Fr-Ci)v^(n-t+1)

5.7963 = ( 1000 * 6% - C * 7% ) * (1/1.07)^(30 -10 + 1) = (60 -0.07 * C ) * 1/1.07^21

5.7963 = (60 -0.07 * C ) * 0.241513

60 - 0.07 *C = 5.7963/0.241513 = 23.9994

0.07 * C = 60-23.9994

C = 514.28